# Pre-trained Encoder Inference: Revealing Upstream Encoders In Downstream Machine Learning Services

### v1

@@ VI-B Results Analysis
Classification accuracies of the $24$ built image classification services ($6$ image encoders $\times$ $4$ downstream image datasets) and the $24$ built text classification services ($6$ text encoders $\times$ $4$ downstream text datasets) are reported as Table A-B2 and Table A-B2 in Appendix A-C. We next show how PEI attacks can reveal encoders hidden in these downstream targets.
@@
The PEI attack is extremely effective against hidden image encoders. The PEI scores and PEI $z$-scores of candidate image encoders in different image classification services are reported in Table I, which shows that our PEI attack succeeds in revealing hidden encoders in all $24$ services. Specifically, in every attack case, the PEI $z$-score of the correct hidden encoder is always the only one that is above the preset threshold of $1.7$ and in most cases, is around $2.0$, which is significantly higher than the preset threshold. This indicates that the PEI attack is very confident in its (correct) predictions.
While the PEI attack is a bit less effective in revealing hidden text encoders, it is strong in avoiding false-positive inferences. For text classifications, the PEI scores and $z$-scores of candidate text encoders are reported in Table II. In all the $24$ downstream cases, the PEI attack correctly infers $14$ hidden encoders, which, however, is less effective than that in image classifications. Further, in some failure cases, although the PEI score of the hidden encoder is the highest among all candidates, its corresponding $z$-score is not above the threshold $1.7$ (e.g., RoBERTa (base) + SST-5) and thus resulted in failure. We deduce this is due to the fact that current synthesized PEI attack texts are gibberish, which makes them difficult to recognize by text encoders pre-trained on meaningful text data and thus could not enjoy semantic embeddings similar to that of the human-readable objective texts (see Table V in Appendix A-B). A potential remedy is to improve the readability of synthesized PEI attack texts.
@@
Nevertheless, it is also worth noting that the PEI attack does not make false-positive inferences at all against text encoders. In every failure case in Table II, the inferred result of the PEI attack is simply $\emptyset$” (which means no significant candidate is found) rather than a wrong candidate encoder. Such a capability of avoiding false-positive inferences is very useful to adversaries, as they would thus not need to make meaningless efforts in tackling false-positive signals. It enables a PEI adversary to try many targets quickly and thus will have a good chance of finding a victim among them in a short time.
The PEI attack seldom makes mistakes when the targeted hidden encoder is not included in the candidate set. We now turn to analyzing the PEI attack in a more challenging setting where the targeted hidden encoder is not in the candidate set. We take the CIFAR-10 dataset for image classifications and the SST-5 dataset for text classifications as examples. In each case, we remove the correct hidden encoder from the candidates set and then launch the PEI attack following Algorithm 1 as usual with the remaining candidates. The $z$-scores of different candidates are presented as Figure 3.
Ideally, since the correct hidden encoder is excluded from the candidate set, all $z$-scores of remaining candidates should not go beyond the preset threshold $1.7$ as none of these candidates are correct. From Figure 3, we find that the PEI attack only makes mistakes in inferring out incorrect candidates in $2$ out of $6$ cases on CIFAR-10 (i.e., RN50 (HF) and MobuleNetV3), and makes all correct inferences (i.e., returning $\emptyset$”) on SST-5. Combined with Table II, all results demonstrate that the PEI attack enjoys a strong and robust capability in avoiding false-positive inferences.
Query budget prices of both image and text encoders can be as low as hundreds of dollars. For image encoders, based on the query budget equation in Section V-A, the exact budget in this part of the experiments is $4$ million per image encoder. According to Table A-B2 in Appendix D, the price of commonly used real-world EaaSs for images is usually around \$$\$0.0001$ per image. Thus, the estimated price of synthesizing PEI attack images would be no more than \$$\$400$ per encoder.
Besides, for text encoders, based on the query budget equation in Section V-B the exact budget in this part of the experiments is no more than $15.4$ million per text encoder. According to Table A-B2 in Appendix D, the price of commonly used real-world EaaSs for texts is at most \$$\$0.014$ per $1,000$ queries or \$$\$0.1$ per $1$ million characters, which thus means the estimated price of synthesizing PEI attack texts in this section is an acceptable price of no more than \$$\$215$ per encoder.
VI-C Ablation Studies
Effect of downstream classifier initialization. The PEI attack in Tables I and II is conducted against single downstream classifiers. Yet, it remains unknown how the PEI attack would be affected by the random initialization of these classifiers. To investigate this, we conduct case studies on CIFAR-10 and SVHN datasets for image classification services and SST-5 and TREC datasets for text classification services. For each dataset-encoder pair, we retrain $5$ downstream models, perform the PEI attack against all of them, and report rates of attack success, false-negative, and false-positive in Figure 4. This leads to a total of $24$ analyzed cases.
From the figure, we have two observations. Firstly, PEI attacks never made false-positive inferences in all analyzed cases. As explained in the previous section, this capability is useful for PEI adversaries to quickly find victims from a large number of targeted services. Secondly, in most of the cases (except three cases “RN50 (MS) + SVHN”, “RoBERTa (base) + TREC”, and “CLIP ViT-L/14 + TREC”), the attack achieved consistent results, i.e., either attack success or false-positive, with a possibility of at least $80\%$, in $21$ out of $24$ analyzed cases. This suggests that the PEI attack is generally insensitive to the initialization of downstream classifiers in image/text classification services.
Figure 4: Rates of different PEI attack results on CIFAR-10, SVHN, SST-5, and TREC datasets. Rates of attack success, false-negative, and false-positive are colored in different colors.
(a) RN34 (HF) + CIFAR-10.
(b) BERT (base) + SST-5.
Figure 5: PEI $z$-scores of candidates on classification services built with downstream classifiers of $4$ different architectures. $z$-scores associated with different services are in different colors.
Effect of downstream classifier architectures. So far, the architecture of downstream classifiers used in image/text classification services is fixed to a $3$-layer MLP with a width of $512$. We now analyze how the architecture of this classifier would affect the PEI attack performance. The analysis is conducted on two cases: “RN34 (HF) + CIFAR-10” and “BERT (base) + SST-5”. For each analyzed case, we build classification services with the targeted encoder and four different downstream classifiers, which are MLP-1, MLP-2, MLP-3, and MLP-4. Here, “MLP-$x$” denotes an MLP architecture of $x$ layers, where dimensions of all hidden layers (if it has) are fixed to $512$. The training of these downstream classifiers follows that described in Section VI-A. After that, we attack each service with PEI attack samples originally used in Section VI-B. PEI $z$-scores are reported in Figure 5.
From the figure, one can find that the downstream classifier architecture has little effect on attacks against vision services (i.e., “RN34 (HF) + CIFAR-10”): the PEI $z$-scores of each candidate are almost the same across different downstream classifier architectures, and the PEI attacks can always reveal the correct hidden encoder. However, for language services (i.e., “BERT (base) + SST-5”), PEI $z$-scores are significantly affected by downstream classifier architectures. In the worst case where the downstream classifier is MLP-1, the PEI attack even failed to infer the correct hidden encoder. Nevertheless, the PEI attack still succeeds in the remaining $3$ cases.
(a) RN34 (HF) + CIFAR-10.
(b) RN50 (MS) + CIFAR-10.
Figure 6: PEI $z$-scores of candidates on image classification services with different gradient estimation sampling number $S$ (see Algorithm 2) varies in $\{25,50,75,100\}$.
Effect of zeroth-order gradient estimation. We now investigate how the zeroth-order gradient estimation in Algorithm 2 would affect the PEI attack against image encoders. Since the gradient estimation accuracy depends on the random sampling number $S$, we thus vary $S$ in the set $\{25,50,75,100\}$ (in Section VI-B, $S=100$) to see how the attack performance would change. This study is conducted on “RN34 (HF) + CIFAR-10” and “RN50 (MS) + CIFAR-10”, and PEI $z$-scores of different encoders under different $S$ are plotted in Figure 6. For the first case (Figure 6(a)), the sampling number $S$ has little effect on PEI $z$-scores, while the $z$-score of the correct hidden encoder always stays in high level (around $2.0$). For the second case (Figure 6(b)), one can find that as the sampling number $S$ increases, the PEI $z$-score of the correct encoder significantly increases, while those of other candidates remain at low levels. These imply that when the PEI attack is weak, it can benefit from a large sampling number $S$. However, one should also notice that a large $S$ would increase the query budget and the adversary needs to carefully trade-off between the utility and efficiency of the PEI attack.
(a) BERT (base) + SST-5.
(b) T5 (base) + SST-5.
Figure 7: PEI $z$-scores on text classification services with different beam-search parameter $K_{1}$ (see Algorithm 3) varies in $\{50,100,150,200\}$.
Effect of beam-search-based optimization. Finally, we analyze how the improved black-box beam search (i.e., Algorithm 3) would affect the PEI attack against text encoders. As the beam search only has two parameters, $K_{1}$ and $K_{2}$, and we usually fix $K_{2}$ to be the half size of the character set $C$, here we only analyze the impact of $K_{1}$ on the PEI attack. The analysis is based on “T5 (base) + SST-5” and “BERT (base) + SST-5”, where the parameter $K_{1}$ varies in the set $\{50,100,150,200\}$ and PEI $z$-scores on different encoders are plotted in Figure 7. Intuitively, a high $K_{1}$ should lead to a strong PEI attack. However, from the figure, we find that as $K_{1}$ increases, the PEI scores of both correct and incorrect encoders change either slightly (Figure 7(a)) or without regular patterns (Figure 7(b)). We deduce this may be partially because current synthesized PEI texts are gibberish, and no matter how well the algorithm is for synthesizing gibberish, they are, in general, difficult to enjoy similar embeddings with those human-readable objective texts. Thus, improving the readability of PEI attack texts may also enhance the PEI attack. We leave this to future studies.
VII Experiments on Text-to-image Generation
This section empirically analyzes the PEI attack against text encoders hidden in real-world text-to-image services.
TABLE III: PEI scores and PEI $z$-scores of candidate encoders on different downstream text-to-image services. If a candidate has $z$-score that is the only one above the threshold, then it is inferred as the encoder hidden in the downstream service.
Downstream Task Candidate Encoder PEI Score (%) / PEI $z$-Score (Threshold = 1.7) Inferred Encoder Attack Success
Model Text Encoder CLIP ViT-B/16 CLIP ViT-B/32 CLIP ViT-L/14 O-CLIP ViT-B/32 O-CLIP ViT-L/14 O-CLIP ViT-H/14
SD-v1.2 CLIP ViT-L/14 0.41 / -0.24 0.39 / -0.64 0.47 / 1.95 0.40 / -0.44 0.39 / -0.73 0.42 / 0.09 CLIP ViT-L/14
SD-v1.4 CLIP ViT-L/14 0.39 / -0.10 0.38 / -0.53 0.45 / 1.98 0.38 / -0.39 0.37 / -0.79 0.39 / -0.17 CLIP ViT-L/14
SD-v2.1 O-CLIP ViT-H/14 0.39 / -0.50 0.37 / -0.93 0.40 / -0.41 0.41 / -0.23 0.43 / 0.16 0.51 / 1.91 O-CLIP ViT-H/14
SD-v2.1 (base) O-CLIP ViT-H/14 0.39 / -0.50 0.37 / -0.91 0.39 / -0.50 0.40 / -0.12 0.41 / 0.11 0.50 / 1.91 O-CLIP ViT-H/14
VII-A Experimental Setup
Pre-trained encoders. We adopt text encoders from six vision-language models to form the PEI candidate set. They are: CLIP ViT-B/16 [2], CLIP ViT-B/32, CLIP ViT-L/14, OpenCLIP ViT-B/32 [62], OpenCLIP ViT-H/14, and OpenCLIP ViT-L/14. All these pre-trained encoders can be downloaded from the Hugging Face Model Repository. See Table A-B2 in Appendix C for details.
Downstream tasks. We directly use stable diffusion models [63] on the Hugging Face Model Repository as the downstream text-to-image generation task in our experiments (see Table A-B2 in Appendix C for download links). They are: Stable Diffusion-v1.2, Stable Diffusion-v1.4, Stable Diffusion-v2.1 (base), and Stable Diffusion-v2.1. The first two adopt the text encoder in CLIP ViT-L/14, while the last two adopt that in OpenCLIP ViT-H/14.
PEI attack samples synthesis. We use the same $K_{1}=20$ objective texts (see Table V in Appendix A-B) as that in text classification experiments. For each pair of objective text and candidate encoder, we synthesize $K_{2}=20$ PEI attack texts via Algorithm 3. For the chosen of hyperparameters, without explicitly stating, the character set $C$ consists of all $26$ lowercase alphabets, all $10$ numbers, and additional characters “!?.,!@#\$%^&*()-=+”, the text length $T$ as $32$, and the beam-search parameters $K_{1}$ and $K_{2}$ are set as $64$ and $32$.
PEI score calculation. We use the image encoder in CLIP ViT-B/16 to evaluate the behavior similarity in the PEI score calculation in Eq. (2). Concretely, for each pair of input objective and PEI attack texts, we first use the targeted text-to-image service to generate $4$ pairs of images and then calculate the averaged cosine similarity of the $4$ pairs of generated images. This leads to the PEI score calculation as
$\zeta_{i}=\sum_{j,k,o=1}^{M_{1},M_{2},4}\frac{\mathrm{Cosine}(\mathrm{CLIP}(g(x^{(obj)}_{j})^{(o)}),\mathrm{CLIP}((g^{(atk)}_{i,j,k})^{(o)}))}{4M_{1}M_{2}},$
where $\mathrm{Cosine}(\cdot,\cdot)$ is the cosine similarity function, $\mathrm{CLIP}$ denotes the image encoder in CLIP ViT-B/16, and $g(x^{(obj)}_{j})^{(o)}$ and $g(x^{(atk)}_{i,j,k})^{(o)}$ denote the $o$-th pair of generated images.
VII-B Results Analysis & Ablation Studies
Attack performance. PEI scores and $z$-scores of candidate text encoders on the four targeted downstream text-to-image models are reported in Table VII, in which we find that the PEI attack succeeds in revealing hidden text encoders in all of the four targeted models. Additionally, in all cases, PEI $z$-scores of the correct candidates are around $1.9$, which is significantly higher than the preset threshold $1.7$. These results demonstrate that the PEI attack is very effective against text encoders in downstream text-to-image services.
@@ VII-B Results Analysis & Ablation Studies
Figure 8: PEI $z$-scores of candidate encoders on different text-to-image services where the hidden encoder is not in the PEI candidates set. No $z$-score go beyond the preset threshold (i.e., $1.7$) means that the attack makes no false-positive inference.
(a) Stable Diffusion v1.4.
(b) Stable Diffusion v2.1.
Figure 9: PEI $z$-scores on text-to-image services with different beam-search parameter $K_{1}$ (Algorithm 3) in $\{32,64,100\}$.
The PEI attack does not make mistakes even when the targeted hidden encoder is excluded from the candidate set. For each of the $4$ analyzed diffusion models, we remove its corresponding text encoder from the PEI candidates set and perform the PEI attack with the remaining candidates. The results are shown in Figure 8, from which we find that on all the $4$ targeted services, no $z$-score of the incorrect remaining candidates is above the preset threshold $1.7$. That means no incorrect candidate is inferred as the hidden encoders, and the PEI attack does not make a false-positive prediction in all analyzed attack scenarios.
@@
Query budget in price. Based on the query budget equation in Secion V-B, the exact budget in this part of the experiments is no more than $26.5$ million per text encoder. According to the pricing data collected in Table A-B2 in Appendix D, the price of commonly used real-world EaaSs for texts is at most \$$\$0.014$ per $1,000$ queries or \$$\$0.1$ per $1$ million characters. This results in an acceptable price of synthesizing PEI attack texts of no more than \$$\$370$ per encoder.
Effect of beam-search-based optimization. Finally, we conduct an ablation study on the impact of the improved black-box beam search (i.e., Algorithm 3) on the PEI attack against text encoders. Similar to that for text classification services (see Section VI-C), here we also focus on analyzing the beam search parameter $K_{1}$ (as $K_{2}$ is always fixed to be around the half size of the character set $C$). Specifically, the analysis is on the Stable Diffusion v1.4 and v2.1, where the parameter $K_{1}$ varies in the set $\{32,64,100\}$. From the figure, we find that keeping the $K_{1}$ to be sufficiently large (e.g., set $K_{1}=64$ in Figure 8(b)) is necessary to ensure the PEI $z$-score of the correct hidden encoder is significantly above the preset threshold. However, further improving $K_{1}$ has little effect on the PEI attack performance but increases the query budget.
VIII Case Study: PEI-assisted Adversarial Attacks Against LLaVA
So far, we have justified the effectiveness of the PEI attack in revealing hidden encoders in downstream services through extensive experiments. In this section, we further show that the PEI attack itself is useful, by conducting a case study on how the PEI attack can assist adversarial attacks against LLaVA [28, 29], one of the most advanced multimodal models. Our results demonstrate there is indeed a need to design general defenses against the PEI attack.
Overview of LLaVA. The case study is based on the multimodal model LLaVA-1.5-13B [28], which is built upon a finetuned version of the Vicuna-1.5-13B language model [64] and the official version of the CLIP ViT-L/14-336px image encoder [2]. It takes images and texts as inputs and outputs texts, and thus can be used for multimodal tasks such as chatting or question-answering. Both of the LLaVA model and its CLIP image encoder can be downloaded from the Hugging Face model repository (see Table A-B2 in Appendix C).
@@ VIII Case Study: PEI-assisted Adversarial Attacks Against LLaVA
TABLE IV: PEI scores and $z$-scores of candidate image encoders on the LLaVA-v1.5-13b. If a candidate has $z$-score that is the only one above the threshold (i.e., 1.7), then it is inferred as the encoder hidden in the downstream service.
Candidate Image Encoder PEI Score PEI $z$-Score (Threshold = $1.7$)
CLIP-ViT-B/16 0.55 -1.50
CLIP-ViT-B/32 0.64 0.22
CLIP-ViT-L/14 0.60 -0.46
CLIP-ViT-L/14-336px 0.72 1.76
OpenCLIP-ViT-B/32 0.60 -0.47
OpenCLIP-ViT-L/14 0.65 0.45
OpenCLIP-ViT-H/14 0.63 0.00

### v2

@@ 4.2 Results analysis
Classification accuracies of the built $18$ image classification services ($6$ encoders $\times$ $3$ downstream datasets) are reported as Table 5 (Appendix B.3). We then study the performance of the PEI attack.
@@
The PEI attack is extremely effective in image classification services. The PEI scores and $z$-scores of candidate vision encoders in different image classification services are reported in Table 1, showing that our PEI attack successfully revealed hidden encoders in $16$ out of $18$ targeted services. Further, in most of the successfully attacked cases, the PEI $z$-score of the correct hidden encoder can be around $2.0$, which is significantly higher than the preset threshold of $1.7$.
@@
Nevertheless, while the PEI attack sometimes may fail, it does not make false-positive inferences at all. In every failure case in Table 1, the inferred result of the PEI attack is simply “ $\emptyset$” (which means no significant candidate is found) rather than a wrong candidate encoder. Such a capability of avoiding false-positive inferences is very useful to adversaries, as they would not need to make meaningless efforts in tackling false-positive signals. It enables a PEI adversary to try many targets quickly and thus will have a good chance of finding a victim among them in a short time.
@@
Figure 2: PEI $z$-scores of candidates on different CIFAR-10 classification services where the correct hidden encoder is not in the PEI candidates set. $z$-scores that are above the threshold $1.7$ are highlighted. Ideally, none of the reported $z$-score should go beyond the preset threshold.
The PEI attack seldom makes mistakes when the correct hidden encoder is not included in the candidate set. We now turn to analyzing the PEI attack in a more challenging setting where the correct hidden encoder is not in the candidate set. We take the CIFAR-10 dataset for image classifications as an example. In each case, we remove the correct hidden encoder from the candidates set and then launch the PEI attack as usual with the remaining candidates. The $z$-scores of different candidates are presented as Figure 2. Ideally, since the correct hidden encoder is excluded from the candidate set, all $z$-scores of remaining candidates should not go beyond the preset threshold $1.7$ as none of these candidates are correct. From Figure 2, we find that the PEI attack only makes mistakes in inferring out incorrect candidates in $1$ out of $6$ cases on CIFAR-10 (i.e., RN34 (HF)).
@@
The query budget price (per encoder) is low. Based on the query budget equation in Section 3.3, the exact budget in this part of experiments is $1$ million per vision encoder. From Table D in Appendix D, the price of commonly used real-world EaaSs for images is not larger than \$$\$0.0001$ per image. Thus, the estimated price of synthesizing PEI attack images would be around \$$\$100$ per encoder.
Ablation studies. We also conduct experiments to analyze how downstream classifiers and PEI hyperparameters would affect the performance of the PEI attack. See Appendix B.4 for details.
4.3 Case study: PEI-assisted model stealing attack
Finally, we show how the PEI attack can facilitate model stealing attacks [63, 51] against downstream image classification services. Given a targeted downstream ML service with only API access, the goal of model stealing is to train a stolen model to replicate the functionality of the targeted service. Thereby, the adversary will first collect a set of surrogate data and use the targeted service to “label” these data. Then, the stolen model will be trained from these surrogate data labeled by the targeted service. When using the PEI attack to facilitate model stealing, the adversary will instead train a downstream classifier based on the hidden encoder inferred by the PEI attack as the stolen model.
As a case study, here we analyze model stealing attacks against image classification services built upon the “RN-50 (HF)” encoder. For the returned response of each query, we consider two settings: (1) “Soft label” setting, where the targeted service will return predicted logits for all classes, and (2) “hard label” setting, where the targeted service will only return the label of top-1 confidence. Besides, we focus on three stolen models: (1) Correct model, which is a downstream MLP classifier trained from the correct hidden encoder “RN-50 (HF)”. (2) Wrong model, which is a downstream MLP classifier trained from an incorrect candidate encoder “CLIP-ViT/L14”. (3) Scratch model, which is a ResNet-18 trained from scratch on the downstream dataset. Among the three stolen models, the correct model is used to simulate the performance of the PEI-assisted model stealing attack.
Experimental setup. We use $50$ thousand images from ImageNet [14] as the surrogate data. This results in a query budget of $50$ thousand. Following jagielski2020high [34], we adopt two metrics to evaluate the model stealing performance: (1) Accuracy, which is the classification accuracy of the stolen model on downstream test data, and (2) Fidelity, which is the rate of downstream test data for which the stolen and target models have the same top-1 prediction. See Appendix B.5 for more experimental details.
Table 2: Model stealing performance of stolen models on different downstream services. “Correct” is the classifier with the correct hidden encoder, “Wrong” is the classifier with the incorrect candidate encoder, and “Scratch” is the ResNet-18 trained from scratch.
Downstream Task Stolen Model Soft Label (%) Hard Label (%)
Accuracy Fidelity Accuracy Fidelity
CIFAR-10 Correct 91.48 98.16 89.38 92.80
Wrong 77.35 73.69 72.89 69.12
Scratch 19.41 19.38 15.84 15.82
SVHN Correct 55.80 62.02 26.61 28.48
Wrong 8.60 8.16 7.38 7.00
Scratch 14.54 14.25 12.14 11.89
Food-101 Correct 62.61 76.35 49.10 54.71
Wrong 29.62 24.14 23.99 20.48
Scratch 2.40 2.49 3.63 3.65
Results analysis. The model stealing results for different types of stolen models are presented in Table 4.3, in which we have two observations. Firstly, for every targeted downstream service, the model stealing performance of the “correct model” is $2\sim 20$ times better than the “scratch model”. Recall that in Section 4.2, our PEI attack revealed the hidden “RN-50 (HF)” encoder in every downstream service. This means that the model stealing adversary can first leverage our PEI attack to reveal the correct hidden encoder, and then exploit this encoder to train a stolen model of better performance. Secondly, the model stealing performance of the “correct model” is also significantly higher than the “wrong model” in every case. Therefore, to launch a strong model stealing attack, adversaries are better not randomly picking an encoder from the candidate set to perform the attack. Instead, they can leverage our PEI attack to reveal the correct hidden encoder before launching model stealing.
5 Experiments on multimodal generative model
This section studies how the PEI attack can reveal vision encoder hidden in the multimodal generative model, LLaVA [40], and show how the revealed encoder can assist adversarial attacks against LLaVA.
5.1 PEI attack against LLaVA model
Setup. The targeted multimodal generative model is LLaVA-1.5-13B [40], which is built upon a finetuned Vicuna-1.5-13B language model [73] and the origional CLIP ViT-L/14-336px vision encoder [54]. It takes images and texts as inputs and outputs texts, and thus can be used for multimodal tasks such as chatting or question-answering. Our PEI attack goal is to reveal the correct vision encoder used by LLaVA-1.5-13B. To this end, we consider a PEI candidates set consists of seven vision encoders, which are: CLIP ViT-B/16 [54], CLIP ViT-B/32, CLIP ViT-L/14, CLIP ViT-L/14-336px, OpenCLIP ViT-B/32 [12], OpenCLIP ViT-H/14, and OpenCLIP ViT-L/14. For the synthesis of PEI attack images, we adopt the same objective images and hyperparameters as that in Section 4.1. For the PEI score calculation, to evaluate the behavior similarity in Eq. (2), we leverage the question-answering capability of LLaVA to directly score the similarity between objective image and PEI attack images (in the range $[0,1]$). See Appendix C.2 for details of PEI score calculation.
Results. The PEI attack results are presented in Table 2(b). From Table 2(a), we find that the correct hidden encoder, i.e., CLIP Vit-L/14-336px, is the only one that has a PEI $z$-score above the preset threshold $1.7$. Further, when the correct hidden encoder is not in the PEI candidates set, Table 2(b) shows that the PEI attack would not infer out a wrong encoder. All these results demonstrate that our PEI attack can effectively reveal hidden encoders in LLaVA while avoiding false-positive predictions.
5.2 Case study: PEI-assisted adversarial attack
@@
Table 3: PEI scores and $z$-scores of candidate vision encoders on the LLaVA-v1.5-13B. If a candidate has $z$-score that is the only one above the threshold $1.7$, then it is inferred as the hidden encoder.
(a) Attack results when the hidden encoder is in the PEI candidates set.
Candidate Encoder PEI Score / $z$-Score
CLIP-ViT-B/16 0.57 / -1.31
CLIP-ViT-B/32 0.66 / 0.62
CLIP-ViT-L/14 0.61 / -0.37
CLIP-ViT-L/14-336px 0.73 / 1.88
OpenCLIP-ViT-B/32 0.62 / -0.23
OpenCLIP-ViT-L/14 0.61 / -0.37
OpenCLIP-ViT-H/14 0.62 / -0.22
(b) Attack results when the hidden encoder is NOT in the PEI candidates set.
Candidate Encoder PEI Score / $z$-Score
CLIP-ViT-B/16 0.57 / -1.62
CLIP-ViT-B/32 0.66 / 1.52
CLIP-ViT-L/14 0.61 / -0.09
OpenCLIP-ViT-B/32 0.62 / 0.13
OpenCLIP-ViT-L/14 0.61 / -0.09
OpenCLIP-ViT-H/14 0.62 / 0.15
Next, we conduct a case study to show how the vision encoder revealed by the PEI attack can assist adversarial attacks against the targeted LLaVA-1.5-13B. We consider an attack that aims to use visually benign adversarial images to induce LLaVA to generate harmful information. These adversarial images can be used to stealthily spread harmful information, as their harmfulness is difficult for humans to detect.
We synthesize such adversarial examples directly based on the hidden encoder revealed by our PEI attack. To launch the attack, we first construct a targeted image that contains predefined harmful plain text content. Then, since the revealed CLIP encoder is open-source, we can synthesize an adversarial image that has embedding similar to that of the targeted image in a white-box manner. This will result in an adversarial image that contains benign mosaics but LLaVA can read out the predefined harmful content from it. See Appendix C.3 for details about the adversarial image synthesis.
Results. Figure 4.3 presents two examples, in which one can find that although the adversarial images only contain visually benign mosaics, LLaVA reads out false medical/health information from PEI-assisted adversarial images. This verifies the effectiveness of the PEI-assisted adversarial attack and further indicates the usefulness and hazards of PEI attack.